what will be the remainder when 16^2+17^2+18^2+19^2 is divided by 70?
Answer by fractalier(2101)

I do not think there is a short cut...merely calculate...thus
16^2+17^2+18^2+19^2 =
256 + 289 + 324 + 361 = 1230
When you divide by 70, your remainder is 40.